Find the tension T in the deltoid muscle in the following fig? If θ = 16 ͦ and the weight of the arm W1 = 68 N and the weight of the hand W2 =45 N
1
Expert's answer
2018-11-01T16:01:09-0400

T=(W1+W2)·sin⁡Θ

T=(68+45)·sin16 ̊=31.2 N